OpenStudy (anonymous):

The mass of Uranus is about 8.7x10(22power)metric tons. The mass of the sun is about 1.998x10(27power) metric tons. About how many times the mass of Uranus is the mass of the sun? Give answer in scientific notation.

OpenStudy (kira_yamato):

\[\frac{29}{666000}\]

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

THE MASS OF THE SUN IS ABOUT ??? TIMES THE MASS OF URANUS???

OpenStudy (anonymous):

\[\frac{\left(1.998*10^{27}\right)}{\left(8.7*10^{22}\right)}=22965.5=2.297*10^4 \]rounded to four significant figures.

OpenStudy (anonymous):

THANKS